Standing waves on the wire are described by the expression y(x,t)=Asin (kx)sinwt. Supposing that the standing wave amplitude is given to be 2.5mm whose wave number(k) is 0.750pi and the angular frequency(w) is 942rad/s. If the left end of the wire is at x=0. At what distance from the left end are the

(1) Nodes of the standing wave?

(2) Antinodes of the standing wave?


1
Expert's answer
2021-04-05T11:07:13-0400

Solution.

"y(x,t)=Asin (kx)sinwt;"

"A=2.5\sdot10^{-3}m;"

"k=0.750\pi (rad/m);"

"\omega= 942rad/s;"

"x=0;"

"1)sin(kx)=0\implies kx=\pi n, n=0,1,2,3..."

"x=\dfrac{\pi n}{k};" "x=0; x=\dfrac{\pi}{0.750\pi}=1.33m; x=\dfrac{2\pi}{0.750\pi}=2.66m;x=\dfrac{3\pi}{0.750\pi}=3.99m.";

"2)sin(kx)=1 or -1 \implies kx=\dfrac{ \pi}{2}+\pi n, n=0,1,2,3..."

"x=\dfrac{\pi}{2k}+\dfrac{\pi n}{k};" "x=\dfrac{\pi}{2\sdot0.750\pi}=0.67m; x=2m; x=3.34m."


Answer: "1)x=0; x=1.33m; x=2.66m;x=3.99m;"

"2)""x=0.67m; x=2m; x=3.34m."